need help answering this question first person to answer with well explained step by step gets the points
thanks

need help answering this question first person to
Answer #1

Suppose at point A reaction is R1

and B = R2

total load = 3xw+1/2x3xw=3x4+1/2x3x4

=12+6=18

so R1+R2 = 18

second equation from taking moment along A

ドー | L/2where L =3

0=wL xL/2+ 1/2wxL (L+l/3)-R2X(2L)

R2 x 6=12x1.5+6(3+1)

R2X6=18+24

R2=3+4=7

from equation 1

R1=18-7

=11